OpenStudy (anonymous):

-1/4y = 16 A. 64 B. 4 C. -4 D. -64

OpenStudy (ranga):

To isolate y, multiply both sides by -4.

OpenStudy (arbershabani97):

-1/4y*4=16*4 -1y=64 /*(-1) y=....

OpenStudy (charlotte123):

-1/4 ----> -0.25 -0.25 * -64 = 16 Thus - y = 64*

OpenStudy (anonymous):

I get it now. thanks guys :)